How many grams of oxygen are required to burn 13.50 grams of acetylene, C2H2


1
Expert's answer
2021-06-08T12:34:04-0400

2C2H2 + 5O2 = 4CO2 + 2H2O

n (O2) = 5/2 x n (C2H2)

n = m / M

M (O2) = 16 g/mol

M (C2H2) = 26.04 g/mol

n (C2H2) = 13.50 / 26.04 = 0.52 mol

n (O2) = 5/2 x 0.52 = 1.3 mol

m (O2) = 1.3 x 16 = 20.8 g